OpenStudy (anonymous):

I have a question on a test and answering this will help me solve the question. - Is 2x the same as 2x^1 ?

OpenStudy (solomonzelman):

when you say "A" (regardless of what A is), to say "A^1" would be the same exact thing.

OpenStudy (anonymous):

so 3xy^3 is the same thing as 3x^1y^3 ?

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \displaystyle 3xy^3=3x^1y^3 }\)

OpenStudy (solomonzelman):

and even \(\large\color{black}{ \displaystyle 3xy^3=(3x)^1y^3 }\)

OpenStudy (anonymous):

ok thx for the help :)

OpenStudy (solomonzelman):

whenever I can :)

OpenStudy (anonymous):

You know how to simplify polynomials, right?

OpenStudy (solomonzelman):

I hope so :)

OpenStudy (solomonzelman):

which polynomial ?

OpenStudy (anonymous):

Ok then, so would \[(3x^1y^3)(2x^3y^1)\] = \[6x^4y^4\] ?

OpenStudy (solomonzelman):

yes, that is correct.

OpenStudy (anonymous):

Ok cool. I'm still getting used to them and 100% good at them yet.

OpenStudy (solomonzelman):

k, .... practice is all that matters. You will get it!